# Universality of Dissipation across Holographic Interfaces

### v1

@@ 6.1 Liouvillian gap
For each value of the Janus parameter $\kappa$, we extract the quasinormal mode with the smallest absolute imaginary part,
@@
which can be identified as the Liouvillian gap of the dual interface CFT, characterizing the relaxation time of the system Žnidarič (2015); Baumgartner and Narnhofer (2008).
@@
We then define the dimensionless relaxation ratio with and without the interface as a function Barad et al. (2025)
$c_{\text{relax}}\;\equiv\;\frac{\Gamma_{\text{Janus}}(\kappa)}{\Gamma_{\text{AdS}_{3}}(\kappa=1)}(\Delta,h,\phi_{0})\,.$ (48)
By construction, $c_{\text{relax}}=1$ in presence of a topological interface, and is 0 when the interface is totally reflective.
@@ 6.2 Universality of dissipation
The independence of the insertion point of the coupling supports our claim that the dissipation across the interface is a genuine interface observable that does not depend on the details of the excitation. Below we fix $\phi_{0}=0$.
@@ 6.3 $c_{\text{relax}}$
Figure 3: $c_{\text{relax}}$ (red dots) versus the BPS-Janus deformation parameter $\kappa$. The green line is the energy transmission coefficient $c_{LR}=1/\kappa^{2}$, and the blue line is the entanglement entropy across the interface $c_{\text{eff}}=1/\kappa$, which almost exactly fits $c_{\text{relax}}$ up to our numerical precision.
Finally, we examine the universal function $c_{\text{relax}}(\kappa)$ on the BPS-Janus interface parameter and compare it with other interface observables like the energy transmission coefficient $c_{LR}$ Baig et al. (2024) and the entanglement entropy of half space $c_{\text{eff}}$ Gutperle and Miller (2016).
$c_{LR}=\frac{1}{\kappa^{2}}<c_{\text{relax}}(\kappa)\;\approx\;\frac{1}{\kappa}=c_{\text{eff}},$ (51)

### v2

@@ 6.1 Liouvillian gap
Consider a conformal interface seperating two CFTs parametrized by interface parameters $\kappa$. In this paper we restrict to the case that $c_{1}=c_{2}=c$. If we introduce dissipation on one side of the interface as in the previous section, we can extract the quasinormal mode with the smallest absolute imaginary part,
@@
which can then be identified as the Liouvillian gap of the dual interface CFT, characterizing the relaxation time of the system Žnidarič (2015); Baumgartner and Narnhofer (2008). Below we identify the parameter $\kappa$ as the super-Janus deformation parameter in (18).
@@
We then define the dimensionless relaxation ratio of the Liouvillian gaps with and without the interface as a function Barad et al. (2025)
$\frac{c_{\text{relax}}}{c}\;\equiv\;\frac{\Gamma_{\text{Janus}}(\kappa)}{\Gamma_{\text{AdS}_{3}}(\kappa=1)}(\Delta,h,\phi_{0})\,.$ (48)
By construction, $c_{\text{relax}}=c$ in presence of a topological interface, and is 0 when the interface is totally reflective.
@@ 6.2 Universality of dissipation
The independence of the insertion point of the coupling originates from the fact that the $m=0$ component in the plane wave basis dominates the BPS equations above, thus suppressing the angular dependence. It supports our claim that the dissipation across the interface is a genuine interface observable that does not depend on the details of the excitation. Below we fix $\phi_{0}=0$.
@@ 6.3 $c_{\text{relax}}$
Figure 3: $c_{\text{relax}}/c$ (red dots) versus the BPS-Janus deformation parameter $\kappa$. The green line is the energy transmission coefficient $c_{LR}/c=1/\kappa^{2}$, and the blue line is the entanglement entropy across the interface $c_{\text{eff}}/c=1/\kappa$, which almost exactly fits $c_{\text{relax}}/c$ up to our numerical precision.
Finally, we examine the universal function $c_{\text{relax}}(\kappa)$ on the BPS-Janus interface parameter and compare it with other central-charge-like interface observables, such as the energy transmission coefficient $c_{LR}$ Baig et al. (2024) and the entanglement entropy of half space $c_{\text{eff}}$ Gutperle and Miller (2016). In super-Janus dual to half-BPS interface between the 2D $\mathcal{N}=(4,4)$ SCFTs, we have
$\frac{c_{LR}}{c}=\frac{1}{\kappa^{2}}<\frac{c_{\text{relax}}}{c}\;\approx\;\frac{1}{\kappa}=\frac{c_{\text{eff}}}{c},$ (51)